Aztec products wishes to evaluate its cash conversion cycle (CCC). Research by one of the firms financial analysis indicates that on average the firm holds items in inventory for 65 days, pays its suppliers 35 days after purchase, and collects its receivable after 55 days. The firms annual sales (all on credit) are about R2.1 billion, it's cost of goods sold represent about 67%of sales, and purchase represent about 40%of cost of good sold. Assume a 365-day year. 3.1 what is Aztec products cash conversion(CCC)? 3.2 if Aztec could shorten its CCC by 5 days, would it be best to reduce the inventory holding period, reduce the receivable collection period or extend the accounts payable period? Why? 3.3 how should the firm manage its inventory, accounts receivable, and accounts payable in order to reduce the length of its cash conversion cycle?
1
Expert's answer
2020-07-29T09:16:56-0400

3.1 CCC = Days of Sales Outstanding + Days of Inventory Outstanding – Days of Payables Outstanding

CCC = DSO + DIO – DPO

CCC = 65 + 35 – 55 = 45

3.2 In this case, it would be better to extend the accounts payable period. It is because this is not only a way of motivating the company's employees, but it is also a manner of freeing up more of the firm's cash flow.

3.3 The company should reduce days sales outstanding and inventory outstanding and at the same time increasing the days of payable outstanding.